OpenStudy (anonymous):

17. solve the following equation. 6(z - 2) + 4 = 10 A. z=2 B. z=3 C. z=-3 D. z=-4 18. solve the following equation. 96 = -6(-4 - 3y) A. y=2 B. y=4 C. y=-2 D. y= -4 19. solve the following equation. -4y + 8 = 4(2y - 2) - 2(-16 + 8y) A. y=-4 B. y=4 C. y= -2 D. y= -6

OpenStudy (kl0723):

#1) 6z-12+4=10 6z=10+12-4 6z=18 6z=18/6 z=3

OpenStudy (anonymous):

ok what about the rest?

OpenStudy (kl0723):

#2) 96= 24 +18y 18y=96-24 18y=72 y=72/18 y=4

OpenStudy (kl0723):

#3) -4y +8= 8y -8+32-16y -4y+16y-8y=-8+32-8 4y=16 y=16/4 y=4

OpenStudy (kl0723):

:)

OpenStudy (kl0723):

You're welcome...